OpenStudy (anonymous):

consider the repeating decimal 0.111 use the formula you learned in this lesson to express the sum as a ratio of integers. please help!! :)

OpenStudy (anonymous):

Do you have the specific formula from your lesson that you'd like me to use? Or would you just want me to solve it the normal way?

OpenStudy (anonymous):

I'm looking through my notes and I don't see a specific formula for it. it'd be nice if you could please :)

OpenStudy (anonymous):

Ahh, just noticed that its a repeating decimal... Well I can tell you that the answer is 1/9.. Not sure about work for that though... Just kind of something I memorized. If it was a finite decimal I could help you out a bit more. Otherwise, all I can say is that if you have a repeating decimal where only 1 number is repeating ex: .2222222222 or .88888888888, your fractional representation of that number is always going to be the number over 9 ex: .111111111 = 1/9 .33333333333 = 3/9 = 1/3 .7777777777 = 7/9

OpenStudy (anonymous):

That makes a lot more sense now, I totally get it more when you gave me the examples thanks so much!!! :)